OpenStudy (loser66):

Evaluate \(\int\int_R e^{-x-y}dxdy\) where R is the region in the first quadrant which \(x\leq y\) Please, help graph the region

OpenStudy (loser66):



OpenStudy (loser66):

It is not hard to find the value of the integral algebraically. However, I would like to know which region is R

OpenStudy (loser66):

To me, region 1 is the correct one.

OpenStudy (kainui):

Wait it says \(x \le y\). So like x=1 and y=2 is \(x\le y\) right?

OpenStudy (loser66):

Yes

OpenStudy (kainui):

I'd say that the point (1,2) is in region 1

OpenStudy (loser66):

hehehe... the same method I test for the region. :)

OpenStudy (kainui):

seems like a pretty good method to me :)

OpenStudy (loser66):

http://www.math.illinois.edu/~ajh/370/408doubleintegralssol.pdf Problem 7.

OpenStudy (sdfgsdfgs):

i say 1 too as \[x \le y\]

OpenStudy (loser66):

@Kainui I have too much free time, so that I study actuary online. I tried to solve the problem without looking at the solution. However, the solution on the page is not what I got. That is why I post it here.

OpenStudy (loser66):

@sdfgsdfgs Thank you, friend.

OpenStudy (loser66):

oh, sorry, problem 9

OpenStudy (kainui):

That picture is wrong, it doesn't match the bounds on the integral, they even have: \[\huge \int_{x=0}^\infty \int_{y=x}^\infty\] Ok the interval y=x to y=infinity is the region you describe, but not the picture drawn, since that's clearly everything above the line y=x like we're saying.

OpenStudy (loser66):

Thank you so much. Ha!! I realize that studying online is not easy at all.

OpenStudy (kainui):

Yeah their actual solution of it, the calculation looks fine though haha weird, I guess everyone just makes mistakes time to time

OpenStudy (loser66):

@Kainui do you have a job?

OpenStudy (kainui):

Nope, but I kinda need one.

OpenStudy (loser66):

your major is engineer, right?

OpenStudy (kainui):

I have a degree in chemistry, sorta interested in physical chemistry / spectroscopy and organic chemistry and I just love math lol.

OpenStudy (kainui):

Oooh I'll apply thanks, if you find more stuff feel free to send 'em my way!

OpenStudy (loser66):

ok, by other mistake, I got this page also. hahaha... None of them fits me. http://www.lockheedmartinjobs.com/

OpenStudy (loser66):

good luck, my friend